OpenStudy (anonymous):

Find the sum of a 10-term geometric sequence when the first term is 3 and the last term is 59,049 and select the correct answer below.

OpenStudy (anonymous):

Find the sum of a 10-term geometric sequence when the first term is 3 and the last term is 59,049 and select the correct answer below. 177,147 88,572<---my answer 88,575 177,144

OpenStudy (welshfella):

first term = a = 3 10th term = ar^9 = 59,049 ar^9 / a = r^9 r^9 = 59049 / 3= 19683 r = ninth root of 19683 use ur calc to find this

OpenStudy (anonymous):

88,575 is what I got :) Am I right?

OpenStudy (anonymous):

@Mehek14

OpenStudy (welshfella):

S10 = 3 * 3^10 - 1 ------- 3 - 1

OpenStudy (anonymous):

88573

OpenStudy (welshfella):

i make it 88572

OpenStudy (anonymous):

177146/2=88572 :) Thanks so much!

OpenStudy (welshfella):

yw